OpenStudy (anonymous):

Integration question: integral (cos(x))^6

zepdrix (zepdrix):

Hmm so we have to do some annoying integration for this one. After we integrate, we'll be left with an integral involving (cos(x))^4. So what you'll notice is that we'll effectively decrease the power by 2 each time we integrate. There is a "Reduction Formula for Cosine" that you can look up in your book probably. But I'm assuming since it's a homework assignment or whatever, that your teacher wants you to do the steps. Lemme see if I can remember how to set this up...

OpenStudy (anonymous):

Thanks again

OpenStudy (anonymous):

I should be allowed to use the half/double angle formulas though.

zepdrix (zepdrix):

\[\huge \int\limits \cos^6 dx \quad = \quad \int\limits \color{blue}{\cos^5x}\cdot \color{orangered}{\cos x dx}\]We'll do Integration-by-parts,\[\huge \color{blue}{u=\cos^5x},\qquad \color{orangered}{dv=\cos x dx}\]

zepdrix (zepdrix):

Mmmm lemme make sure I didn't make a mistake real quick :D

OpenStudy (anonymous):

Seems correct

OpenStudy (anonymous):

wait, isn't du -5cos^4(x)*sin(x)?

zepdrix (zepdrix):

Ah ty that's where i got messed up :O knew there was suppose to be another power somewhere lol

OpenStudy (anonymous):

:)

zepdrix (zepdrix):

\[\huge \color{blue}{du=-5\sin x\cdot \cos^4 x dx}, \qquad \color{orangered}{v=\sin x}\] \[\large =\color{blue}{\cos^5x}\color{orangered}{\sin x}+\color{blue}{5}\int\limits\limits \color{blue}{\sin x\cdot \cos^4 x}\color{orangered}{sinx}\color{blue}{dx}\]

zepdrix (zepdrix):

Ignoring the first part for a moment, we have,\[\large 5\int\limits \cos^4x \sin^2x dx \quad = \quad 5\int\limits \cos^4x dx-5 \int\limits \cos^6x dx\]

zepdrix (zepdrix):

I skipped a couple steps in there, let me know if it's too confusing. I changed sin^2 to (1-cos^2), then distributed the cos^4, and wrote it as 2 separate integrals.

OpenStudy (anonymous):

yeah its cool

zepdrix (zepdrix):

Similar to that annoying problem we did yesterday :) We end up with the same thing we started with!\[\large Y=\cos^5x \sin x+5\int\limits \cos^4x dx-5Y\]

zepdrix (zepdrix):

\[\large 6Y=\cos^5x \sin x+5\int\limits\limits \cos^4x dx\] \[\large Y=\frac{1}{6}\cos^5x \sin x+\frac{5}{6}\int\limits\limits \cos^4x dx\]

zepdrix (zepdrix):

From this point, you might want to try and recognize the pattern. And derive the reduction formula from what we're doing. Otherwise doing this 2 more times will be a BEAST!!

OpenStudy (anonymous):

lol, Thanks again :)

zepdrix (zepdrix):

\[\large \int\limits \cos^n dx=\frac{1}{n}\cos^{n-1}x \sin x+\frac{n-1}{n}\int\limits \cos^{n-2}x dx\] This is the reduction formula, if you compare it to what we have so far, it will match up.

zepdrix (zepdrix):

When you do the next reduction, make sure you multiply all the pieces by that 5/6 fraction. Then for the final cos^2, I think you had the right idea earlier, Half Angle! :)

zepdrix (zepdrix):

all the new pieces*

OpenStudy (anonymous):

Thanks I'm going to try that on my test tomorrow :)

OpenStudy (anonymous):

Why don't we learn these in class -_-

zepdrix (zepdrix):

Hmm good question :D It's one of those long tedious problems, so it's less likely to show up on a test. But you know your teacher better than I do :) So I dunno.

OpenStudy (anonymous):

yeah, thanks!